Please show all work

Yield to maturity. (a) Find the yield to maturity on a 1-month T-bill with $100,000 face value that sold for a price of $99,667.77.

Answer: i = 0.04

(b) Consider a loan of $400,000 to be repaid with annual payments over 25 years. Find the fixed payments that would be required for the yield to maturity on the loan to be 6%.

Answer: FP = $31,290.68.

(c) Consider a loan of $200,000 to be repaid with annual payments over 30 years. If the annual fixed payments are $19,467.28, then what is the yield to maturity on the loan? Show your work.

Answer: i = 0.09
Answer #1

a) YTM of 1-month T-bill with 100000 =(Par Value/Face Value)^n-1 =(100000/99667.77)^12 -1 =0.0407

b) PV =400000
Number of Years =25
Rate =6%
Fixed Payment =PV/(1-(1+r)^-n)/r =400000/(1-(1+6%)^-25)/6% =31290.69

c) PV =200000
Fixed Payment =19367.28
Number of Periods =30
YTM using Rate function of excel  =RATE(30,19367.28,-200000) =0.0894 or 0.09